1.Find all the asymptotes for the graph of the rational function f(x) = 3x^2 -1/ x+3
2. Find all the asymptotes of the graph of the rational function f(x) = 4x^2 + 7/ 3(x+1)(x-2)
3. Find all the asymptotes of the graph of the rational function f(x) = 3/x^2

Answer by stanbon(75887)

1.Find all the asymptotes for the graph of the rational function
f(x) = 3x^2 -1/ x+3
----------
Vertical at x=-3
Horizontal aty = 3x^2/0x^2 = 3/0; so no horizontal
--------------------------
2. Find all the asymptotes of the graph of the rational function
f(x) = (4x^2 + 7)/ (3(x+1)(x-2))
Vertical at x = -1 and x=2
Horizontal at y = 4x^2/3x^2 = 4/3
------------------------------------------

3. Find all the asymptotes of the graph of the rational function
f(x) = 3/x^2
Vertical at x = 0
Horizontal at y = 0x^2/x^2 = 0/1 = 0
=========================================
Cheers,
Stan H.